Journalist Report – February 23th

Sol 11 – The sweetness of success

« “How do you like Mars, Pop?”

“Fine. Always something new. I made up my mind when I came here last year I wouldn’t expect nothing, nor ask nothing, nor be surprised at nothing.” »

– Chapter 11 of The Martian Chronicles by Ray Bradbury

This morning, we woke up expecting only to be surprised… and we were!

The forecast announced slightly less fierce winds, allowing the third photogrammetry EVA at North Ridge to take place. Beforehand, the exploration team was shown a 3D render of the same area, and given indications on where to find the checkpoints. The experiment was a success! Jérémy and Alexandre found all six checkpoints much faster and more efficiently than the team with the classic 2D map, and had time left on their EVA to explore North Ridge further. They even mentioned that having seen the 3D map, they felt like they had already been to North Ridge before even going on EVA. Alice was HabCom for the first time, and had fun jotting down notes for the EVA crew to read upon their return.

Next week, the same process will be repeated at a different location. Alice and Quentin are aiming higher: more checkpoints to find, a greater search radius, and more complex terrain.

The conditions in which we have been conducting the EchoFinder experiment have also improved! Thanks to Quentin and the researchers from MEDES, we have been able to transfer the software to a more powerful tablet, making it easier to detect the QR cubes and use the Augmented Reality interface. After Quentin and I performed a test run, Adrien and Alexandre were able to run a full ultrasound session which was successful. Meanwhile, I went to battle with my computer to start creating a mid-rotation video using footage from a first-person camera I wore during EVAs and from our Parrot drones. I cannot believe we are nearing mid-rotation… time flies!

The Hab smelled of cake and caramel all day, doing wonders for the crew’s morale: using up the last of our flour and dough mixes, Corentin and Adrien baked a chocolate chip cake and a brioche!

EVA Report – February 23th

Crew 275 EVA Report 23-02-2023

EVA # 10

Author: Jeremy Rabineau (Commander and EVA Leader)

Purpose of EVA: To perform the photogrammetry experiment with the 3D map.

We went to North Ridge, where the two subjects of the experiment tried to find some specific markers while the operator of the experiment onsite took time measurements and drew their trajectory on a map. In addition, the two subjects were equipped with a step counter.

Start time: 09:08

End time: 10:55


09:08 Airlock closed.

09:13 Exiting the airlock.

09:17 The EVA Crew leaves the Hab with the rovers (Spirit and Curiosity).

09:26 The EVA Crew arrives at North Ridge. EVA crew cannot hear EVA leader when he is on channel 1, so he switches to channel 2.

09:37 The EVA Crew has already reached the two first checkpoints.

09:49 The EVA Crew has reached the five first checkpoints.

09:54 Mission accomplished: The EVA Crew has successfully reached all the checkpoints. They will take some time for pictures and walking a bit around North Ridge.

10:22 The EVA Crew decides to go back to the rovers.

10:27 The EVA Crew is at the rovers and starts its way back to the Hab.

10:36 Both rovers are safely parked at the MDRS.

10:38 Airlock closed.

Destination: North Ridge.

Coordinates (use UTM WGS 84): N4252000 E519000

Participants: EVA Leader: Jeremy Rabineau (Crew Commander & Subject), Alexandre Vinas (Crew Astronomer & Subject), Quentin Royer (Crew Engineer & Operator)

Hab Com: Alice Chapiron (Crew Scientist)

Road(s) and routes per MDRS Map: Cow Dung Road until North Ridge, near the intersection with “Gateway to Candor”, we will then walk towards North-West for an estimated 600 meters from the rover. (See map attached)

Mode of travel: Driving and walking.

GreenHab Report – February 23th

GreenHab Officer: Adrien Tison

Environmental control: Heater

Average Temperature: 74.5 °F

Average humidity: 27 %

Hours of supplemental light: 0 hour

Daily water usage for crops: 8.98 gallons

Daily water usage for research and/or other purposes: 2.12 gallons

Water in the blue tank: 64.9 gallons

Time(s) of watering for crops: 13:34

Changes to crops: N/A

Narrative: Nothing special to signal today, it was a pretty common day.

I maintained and watered plants and crops as my daily operations and checked everything to be sure everything is doing well. In the next few days there will be a really nice amount of tomatoes, they are soon to be harvested!

I also took care of my Microgreens experiment and added water to the fish tank as it has lowered a bit since the beginning.

Harvest: N/A

Support/supplies needed: N/A

Operations Report – February 23th

Crew 275 Operations Report 23Feb2023
SOL: 11

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet
• Spacesuit n°9

Notes on non-nominal systems: None

Spirit rover used: Yes
• Hours: 209.4
• Beginning charge: 100%
• Ending charge: 70 %
• Currently charging: Yes

Opportunity rover used: No
• Hours: 113.2
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Curiosity rover used: Yes
• Hours: 221.2
• Beginning charge: 100%
• Ending charge: 73%
• Currently charging: Yes

Perseverance rover used: No
• Hours: 255.6
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

General notes on rovers: Rovers were used during today’s EVA. Nothing to report.

Summary of Hab operations: The Crew conducted routine operations.
• WATER USE: 12.36 gallons
• Water: 316.72 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: Yes

Summary of internet: Hughes Net was used during the day before the opening of COMMS for email purposes, and Starlink (MDRSCampus-Guest) is used during COMMS. MDRScampus is used for experimental purpose.

Summary of suits and radios: Suits and radios were used during today’s EVA. Nothing to report.

Summary of GreenHab operations: The Crew botanist conducted routine operations.
• WATER USE: 11.10 gallons
• Heater: On
• Supplemental light: No
• Harvest: None

Summary of ScienceDome operations: The Crew Scientist conducted routine operations, and tests for the Human Factors experiment were performed.
• Dual split: Switched off during the night

Summary of RAM operations: No operations were conducted in the RAM today.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns, and requests to Mission Support:
• Do you prefer that we put Spacesuit #9 in the RAM or in the Engineering Airlock? It will be placed at your preferred location tonight.

Sol Summary Report – February 23th

Crew 275 Sol Summary Report 23Feb2023

Sol: 11

Summary Title: The sweetness of success

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Today, the EVA Crew going to North Ridge had the opportunity to study a 3D map of the area before starting their exploration tasks. They reported that they felt almost like they already knew the place and managed to reach all the targeted checkpoints very quickly. This seems to show that such 3D maps generated by drones are very effective in training crews to explore a place they have never been to before. It was actually so effective that they came back about an hour earlier than planned. We decided to spend this spare time cooking, and ended up with a very tasty brioche and an amazing cake!

Inside the Hab, we continued our activities, including installing the software dedicated to AI-assisted echography on a more powerful tablet. It is now even easier for non-trained subjects like us to perform ultrasounds on each other.

In the context of our experiment with the Royal Institute of Technology (Sweden), our Crew Engineer is now working on a location tracking system that will enable the possibility to correlate our stress and emotions to our position in the station.

Look Ahead Plan:

If the weather conditions allow it, we would like to perform the tasks planned for EVA #9 (which was cancelled) tomorrow morning (Sol 12). We would thus re-install the antenna of Mega-Ares and the weather instruments that have been removed during EVA #7. As no rain is forecasted for the next few days, we will also re-install the LOAC.

On Saturday (Sol 13), we would like to perform our first EVA dedicated to the MetMet experiment. The plan would be to collect geological samples based on the information given by a user-friendly instrument measuring the magnetic susceptibility and conductivity of rocks (the aim being to check the interest of a rock sample before collecting it). In addition to this, we would change the batteries of the atmospheric instruments, because we do not plan any EVA for Sunday (Sol 14) and the batteries cannot last more than 48 hours in cold weather.

Anomalies in work: None.

Weather: Windy,low -6°C/22°F, high 3°C/37°F.

Crew Physical Status: Optimal.

EVA: EVA #10 at North Ridge went smoothly, and the EVA Crew managed to reach all the desired checkpoints very quickly.

Reports to be filed: Sol Summary, EVA Request, EVA Report, Operations Report, Journalist Report, GreenHab Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site